IR-IR DOUBLE RESONANCE EXPERIMENT OF CH_{3}F-(ortho-H_{2})_{n} CLUSTERS IN SOLID para-H_{2}

Y. MIYAMOTO, Graduate School of Natural Science and Technology, Okayama University, 3-1-1 Tsushima-naka Okayama 700-8530 JAPAN; A. MIZOGUCHI, and \underline{H. KANAMORI}, Department of Physics, Tokyo Institute of Technology, Tokyo, 152-8551 JAPAN.

FTIR absorption spectrum of the ν_{3} vibrational band of CH_{3}F in solid para-H_{2} shows a series of lines at regular intervals, which could be assigned to clusters forming CH_{3}F-(ortho-H_{2})_{n}, with n = 0 to 12^{a}. However, such a distinct feature appears only in the ν_{3} mode but is not observed in the other vibrational modes. In order to solve this problem, we applied IR-IR double resonance spectroscopy using 3 μm OPO laser and 9 μm QC laser to the vibrational bands of CH_{3}F in solid para-H_{2}. Because, very recently, high resolution cw-QC laser spectroscopy demonstrated a perfect bleaching effect on a spectral peak in ν_{3} mode^{b}. By bleaching the n=1 component of the ν_{3} band at 9 μm, we can see a corresponding depletion in the broad spectral feature of the ν_{1} band at 3 μm. This simultaneous disappearance of the spectral peaks is a proof of the CH_{3}F-(ortho-H_{2})_{n} cluster model, and it suggests that all the vibrational bands should be explained by the same model.

aK. Yoshioka and D. T. Anderson, J. Chem. Phys. \underline{119}, 4731(2003)

bA. R. W. McKellar, A. Mizoguchi, and H. Kanamori, Phys. Chem. Chem. Phys., \underline{13}, 11587 (2011)